OpenStudy (anonymous):

How do you solve 4(x-1)=(x+2)+6

OpenStudy (anonymous):

expand out the left side, then expand out the right side then solve for x

OpenStudy (anonymous):

4x-4=x+8

OpenStudy (anonymous):

You're going to distribute the 4 on the left side 4(x-1) which will equal 4x-4 and the right side you will combine like terms, which will be x+8. So now you have 4x-4=x+8. Now you're going to solve for x.

OpenStudy (anonymous):

Expand the left side. 4(x-1) --> becomes 4x-4. Expand right side. (x+2)+6 --> becomes x + 2 + 6. you can take out parentheses because there is nothing to do inside the parentheses. You end up with x+8. Now you have 4x-4 = x+8 Bring all of the x's to one side... 3x-4 = 8. See the difference? Now bring your constants to the other side (right side) 3x=12... and divide the left by right... x=4